Jack and Jim start at the same time to ride from place A to place B, 60 miles away. Jack travels 4miles an hour slower than Jim. Jim reaches B and at once turns back meeting Jack 12 miles from B. The rate of Jack was:
(A)4 mph (B) 8 mph (C) 12 mph (D) 16 mph (E) 20 mph

Answer by checkley77(12844)

D=RT
72=RT OR T=72/R JIM'S EQUATION.
------------------
60-12=(R-4)T
48=(R-4)T
T=48/(R-4)
72/R=48(R-4)
48R=72(R-4)
48R=72R-288
48R-72R=488
24R=288
R=288/24
R=12 MPH. WAS JIM'S RATE.
12-4=8 MPH. WAS JACK'S RATE. ANSWER. B)
PROOF:
72/12=48/8
6=6